Financial Statement Notes: Quarterly Data
Quarterly data are presented below for Company A and Company B. One of these companies is Gibson Greetings, Inc., which manufactures and sells greeting cards. The other company is Hon Industries, Inc., which manufactures and sells office furniture. Both companies are on a calendar year basis.

(Amounts in Thousands)
First
Quarter
Second
Quarter
Third
Quarter
Fourth
Quarter

Year
Company A
Net sales $186,111 $177,537 $203,070 $213,608 $780,326
Gross profit 55,457 53,643 64,024 69,374 242,498
Company B
Net sales $84,896 $83,796 $142,137 $235,336 $546,165
Gross profit 53,900 52,983 66,018 104,961 277,862


Required
a. Compute the percent of annual net sales generated each quarter by Company A. Round to the nearest percent.
b. Compute the percent of annual net sales generated each quarter by Company B. Round to the nearest percent.

First
Quarter
Second
Quarter
Third
Quarter
Fourth
Quarter
a. Company A annual net sales percent
b. Company B annual net sales percent


c. Which company has the most seasonal business?

Company A or Company B

d. Which Company is Gibson Greetings?

Company A or Company B

Which Company is Hon Industries, Inc.?

Company A or Company B

e. Which company's interim quarterly data are probably most useful for predicting annual results?

Company A or Company B
Answer #1
a
First Quarter Second Quarter Third Quarter Fourth Quarter
a. Company A annual net sales percent 24% 23% 26% 27%
b. Company B annual net sales percent 16% 15% 26% 43%
b
Company B has the most seasonal business.
c
Company B is Gibson Greetings
d
Company A is Hon Industries, Inc.
e
Company A quarterly data is most useful for predicting annual results
Workings for a and b
First Quarter Second Quarter Third Quarter Fourth Quarter
a. Company A annual net sales percent =186111/780326 =177537/780326 =203070/780326 =213608/780326
b. Company B annual net sales percent =84896/546165 =83796/546165 =142137/546165 =235336/546165